(1 point) Let 4 4 -4 3 -2 A-14-3-2 4 -2 5 3 5 Give a non-zero vector in the null space of A
Answer #1

A=\begin{pmatrix}-4&4&-4&3&-2\\ 4&-3&-2&2&2\\ 4&-2&5&3&-5\end{pmatrix}

-4 4 -4 3 -2
4 -3 -2 2 2
4 -2 5 3 -5

.

convert into Reduced Row Eschelon Form...

Divide row1 by -4

1 -1 1 -3/4 1/2
4 -3 -2 2 2
4 -2 5 3 -5


Add (-4 * row1) to row2

1 -1 1 -3/4 1/2
0 1 -6 5 0
4 -2 5 3 -5


Add (-4 * row1) to row3

1 -1 1 -3/4 1/2
0 1 -6 5 0
0 2 1 6 -7


Add (-2 * row2) to row3

1 -1 1 -3/4 1/2
0 1 -6 5 0
0 0 13 -4 -7


Divide row3 by 13

1 -1 1 -3/4 1/2
0 1 -6 5 0
0 0 1 -4/13 -7/13


Add (6 * row3) to row2

1 -1 1 -3/4 1/2
0 1 0 41/13 -42/13
0 0 1 -4/13 -7/13


Add (-1 * row3) to row1

1 -1 0 -23/52 27/26
0 1 0 41/13 -42/13
0 0 1 -4/13 -7/13


Add (1 * row2) to row1

1 0 0 141/52 -57/26
0 1 0 41/13 -42/13
0 0 1 -4/13 -7/13

reduced system is

\begin{pmatrix}1&0&0&\frac{141}{52}&-\frac{57}{26}\\ \:0&1&0&\frac{41}{13}&-\frac{42}{13}\\ \:0&0&1&-\frac{4}{13}&-\frac{7}{13}\end{pmatrix}\begin{pmatrix}a\\ b\\ c\\ d\\ e\end{pmatrix}=0

a+\frac{141}{52}d-\frac{57}{26}e=0......................a=-\frac{141}{52}d+\frac{57}{26}e

b+\frac{41}{13}d-\frac{42}{13}e=0...........................b=-\frac{41}{13}d+\frac{42}{13}e

c-\frac{4}{13}d-\frac{7}{13}e=0.........................c=\frac{4}{13}d+\frac{7}{13}e

d=d

e=e

.

general solution is

\begin{pmatrix}a\\ b\\ c\\ d\\ e\end{pmatrix}=d\begin{pmatrix}-\frac{141}{52}\\ -\frac{41}{13}\\ \frac{4}{13}\\ 1\\ 0\end{pmatrix}+e\begin{pmatrix}\frac{57}{26}\\ \frac{42}{13}\\ \frac{7}{13}\\ 0\\ 1\end{pmatrix}

.

null space are

{\color{Red} \begin{pmatrix}-\frac{141}{52}\\ -\frac{41}{13}\\ \frac{4}{13}\\ 1\\ 0\end{pmatrix},\:\begin{pmatrix}\frac{57}{26}\\ \frac{42}{13}\\ \frac{7}{13}\\ 0\\ 1\end{pmatrix}}